5. c) Find the total energy of a moving object of mass 6.3 kg and momentum 11.5*10^ 6 kgms^ -1 .

1
Expert's answer
2021-12-09T09:23:29-0500
"E=\frac{mv^2}2=\frac{m(p/m)^2}{2}=\frac{p^2}{2m}=1.05·10^{13}\text{ J}."